OpenStudy (anonymous):

what changes we make in different quadrants to find argument..give the difference 1)Convert the given complex number in polar form: 1 – i 2)Convert the given complex number in polar form: – 1 + i

OpenStudy (anonymous):

@ganeshie8 help me plzz sir i have a test tomorrow

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

\[1-\iota=r \left( \cos \theta+\iota \sin \theta \right)=r \cos \theta +\iota r \sin \theta\] \[r \cos \theta=1,r \sin \theta=-1\] square and add \[r^2\left( \cos ^2\theta +\sin ^2 \theta \right)=1^2+\left( -1 \right)^2=2,r=\sqrt{2}\] \[\sqrt{2}\sin \theta=-1,\sin \theta=-\frac{ 1 }{ \sqrt{2} }=-\sin \frac{ \pi }{ 4 }=\sin \left( \frac{ - \pi }{ 4 } \right)\] \[=\sin \left( 2 \pi -\frac{ \pi }{ 4 } \right)=\sin \frac{ 7 \pi }{ 4 },\theta=\frac{ 7 \pi }{ 4 }\] \[1-\iota =\sqrt{2}\cos \frac{ 7 \pi }{ 4 }+\iota \sqrt{2}\sin \frac{ 7 \pi }{ 4 }\]

OpenStudy (anonymous):